OpenStudy (anonymous):

Consider the infinite geometric series a. Write the first four terms of the series. b. Does the series diverge or converge? c. If the series has a sum, find the sum.

OpenStudy (anonymous):

\sum_{n=-1}^{\infty} -4(1/3)^{n-1}

OpenStudy (anonymous):

is this right for a. ? 1. -4(1/3)^(1-1) = -4 x 1 =-4 2. -4(1/3)^(2-1) = -4/3 3. -4(1/3)^(3-1) = -4/9 4. -4(1/3)^(4-1) = -4/27

OpenStudy (anonymous):

b. ? The series converges.

OpenStudy (anonymous):

and is this correct for c because i have a feeling its wrong? sum= -4/(1-(1/3))=-4/(2/3)=-6